# LaoBench: A Large-Scale Multidimensional Lao Benchmark for Large Language Models

### v2

@@ 3.2 Closed-Form Evaluation on Lao-7k
Figure 5: Radar chart comparing performance of models with and without Chain-of-Thought (CoT) prompting on Lao-7k. Models with CoT (Thinking) consistently outperform their Non-Thinking counterparts, especially in complex reasoning subdomains such as Thinking & Philosophy and Knowledge Application.
@@ 4 Related Work
Multilingual benchmarks such as FLORES Goyal et al. (2021) and large-scale MT evaluation datasets provide broad language coverage, but they often focus on translation or surface-level understanding and typically exclude Lao or provide limited Lao-specific evaluation. Moreover, multilingual benchmarks rarely include culturally grounded knowledge reasoning or education-aligned tasks that reflect real-world language use in Laos.
@@
High-quality benchmarks increasingly combine human expertise with automated verification to ensure scalability and reliability Hendrycks et al. (2021); Luo et al. (2024); Wang et al. (2025); Kang et al. (2025). In particular, recent work highlights the importance of held-out test sets and black-box evaluation to mitigate contamination Deng et al. (2024). LaoBench follows this trend by integrating expert curation with agent-assisted validation and providing a large closed-source subset.

### v3

@@ 3.2 Closed-Form Evaluation on Lao-7k
Figure 5: Radar chart comparing performance of models with and without Chain-of-Thought (CoT) prompting on Lao-7k.
@@ 4 Related Work
Multilingual benchmarks and large-scale MT evaluation datasets provide broad language coverage, but they often focus on translation or surface-level understanding and typically exclude Lao or provide limited Lao-specific evaluation. Moreover, multilingual benchmarks rarely include culturally grounded knowledge reasoning or education-aligned tasks that reflect real-world language use in Laos.
@@
High-quality benchmarks increasingly combine human expertise with automated verification to ensure scalability and reliability Hendrycks et al. (2021); Luo et al. (2024); Wang et al. (2025); Kang et al. (2025); Feng et al. (2025); Zheng et al. (2026); Kang et al. (2026); Feng et al. (2025). In particular, recent work highlights the importance of held-out test sets and black-box evaluation to mitigate contamination Deng et al. (2024). LaoBench follows this trend by integrating expert curation with agent-assisted validation and providing a large closed-source subset.
@@ B.6 Inter-annotator agreement and adjudication
B.7 Expert Contributor Profiles and Transparency
To ensure the linguistic integrity and cultural authenticity of LaoBench, we engaged a diverse group of 55 human contributors throughout the construction, validation, and auditing phases. Their roles, academic backgrounds, and specific responsibilities are summarized in Table 6.
All culturally grounded and curriculum-aligned items underwent a double-blind review process by at least two qualified contributors. Items identified with semantic ambiguity, insufficient contextual grounding, or potential cultural insensitivity were either significantly revised or discarded. Final dataset audits were conducted by senior expert reviewers (predominantly PhD-level faculty) to guarantee alignment with Lao national educational standards.
Role Background Task Description Education Level (Count) Total
Domain Experts Lao universities & educational institutions Culturally grounded knowledge & K12 dataset creation Bachelor (16), Master (5), PhD (4) 25
Linguistic Experts Lao–Chinese / Lao–English translation scholars Parallel corpus construction & random spot-check auditing Bachelor (6), Master (3), PhD (2) 11
Senior Reviewers Lao universities & Dept. of Education Final comprehensive full audit across all categories Master (2), PhD (8) 10
Data Curators Computer Science & NLP backgrounds Data consolidation, formatting, & consistency verification Bachelor (5), Master (2), PhD (2) 9
Total 55
Table 6: Summary of human contributors involved in the LaoBench pipeline. The team composition ensures a balance between pedagogical expertise, linguistic precision, and technical data integrity.